OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = –1.7x + 8.5 Options: 17x + 10y = 85 –17x + 10y = 85 –17x + 10y = –85 17x + 10y = –85

OpenStudy (anonymous):

Sorry I know I've been asking a lot of questions I'm just going through a workpage and it's not in a test or anything I'm just doing the optional questions to try and learn this stuff before my test.

OpenStudy (e.mccormick):

"Use integer coefficients. " know what that means?

OpenStudy (anonymous):

Not really sure in the context that they are asking at least.

OpenStudy (e.mccormick):

Well, integers are the numbers from \(-\infty\) to \(\infty\) that are not fractions, decimals, etc. So 1, 3, -8, and even \(\tfrac{2}{2}\) because it is really 1. A coefficient is a number that is next to a variable. So in 5x, 5 is the coefficent and x is the variable.

OpenStudy (anonymous):

Ok, I get that all. I just don't get how to apply it to the problem.

OpenStudy (e.mccormick):

That means the place to start is to change \(y = –1.7x + 8.5\) so that is has no decimals.

OpenStudy (e.mccormick):

Just remember, if you multiply the right hand side of the equation by something, you must also multiply the left by the same something!

OpenStudy (anonymous):

I'm confused as to how to do this...

OpenStudy (e.mccormick):

Well, if I have 1.1 and multiply by 10, what do I get?

OpenStudy (anonymous):

11

OpenStudy (e.mccormick):

OK. And if I have \(y = –1.7x + 8.5\) and I multiply through the entire equation by 10, what do I get?

OpenStudy (anonymous):

-1.7 = -17 and 8.5 = 85, so would it be 10y = -17x + 85?

OpenStudy (e.mccormick):

Correct. Now, \(10y = –17x + 85\) does not quite look like any of your answers. The \(-17x\) is still in the wrong place.

OpenStudy (anonymous):

So in standard form would it be -17x + 10y = 85

OpenStudy (e.mccormick):

Not quite... to move domething from left to right you add, subtract, multiply, or divide. But, if it is subtracted, like \(-17x\) is, you have to ADD it to move it.

OpenStudy (anonymous):

Oh so it would be 17 not -17

OpenStudy (e.mccormick):

=)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (e.mccormick):

np. I hope you begin to see how these work. It is all about order of operations!